OpenStudy (anonymous):

@LynFran @DecentNabeel @ganeshie8 @jtvatsim @Astrophysics

OpenStudy (anonymous):

I think that it's none of the above

OpenStudy (anonymous):

@LynFran @DecentNabeel @ganeshie8 @jtvatsim @Astrophysics

OpenStudy (anonymous):

btw im here even though it says im not

OpenStudy (lynfran):

we solve the denominator to see where its discontinues/undefined/excluded so b-2=0 b-2+2=0+2 b=?

OpenStudy (anonymous):

2?

OpenStudy (lynfran):

correct